Prev Next

SW&IT Meeting Minutes 2018-04-02

Attendees: Andrew, David, Doug, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Previous Action Items

  • NetBrain (David - schedule demo)
  • VM/SSD (David - followup meeting)
  • binoccular planning follow-up (Joe, Doug, Matthieu - schedule demo) DONE
  • Core A (Stephen - swap during eng night) DONE
  • Argos (Andrew - s/w handover doc progress, schedule meeting) DONE
  • Logging (Yang - requirements, design, concepts iteration - schedule meeting) - TCS lib (PK), other info in the Twiki (TBD)
  • SSP final draft (Matthieu) DONE
  • techtalk#1 (Matthieu) PENDING
  • github user details, responsibilities, etc. (--- - send to MB) DONE
  • file vacation and time off in time DONE

Outlook

Team reports

AC

Last week
  • ARGOS handover
This week
  • ARGOS handover
  • Investigate ITs

DE

DS

Last Week:
  • Pointing & Collimation E time prep
  • Binocular Planning discussions/emails
This Week:
  • Power outage meeting support
  • Pointing & Collimation E time post-mortem & improvement actions

IS

This week:
  • Working on any OT issues
  • but mostly working on two LUCI issuetracks:
    • #6138 by adding a summary fixed row on the LUCI scheduler
    • #7122 this although has a medium priority is really a low priority. The interest on this issue is to understand how LUCI interacts with the MASK robot unit a source of contention and problems in the instrument. By knowing the code a bit better, I hope to be able to add some simply simulation to it. The problem reported happens or originates in what I think is the non-simulated part.

JA

Last week -

Tested the Planner Copointing with MB,PK,DS (remote).
  • Need to clean up the Absorb case. DONE
  • Need to add the contribution of offsets into the final location of the mount position.
This week -

Planner
  • Indicate out-of-copoint steps on timeline.
  • Provide a means to hide all preset header info on Timeline (DS). DONE.
  • Produce some additional test cases. Single offsets.
LUCI mask change
  • New 0.5" x 220" LUCI Longslit mask proposed to replace existing 0.5" x 233" mask
  • Changed is trivial but will require OT rebuild. Just say when.
GitHub
  • The OT + Planner will transition to github.
  • Planner will be a dev branch until release date.
I'm on Software Support duty this month.
  • With the exception of the last four days (KS time swap).
  • Fixed some layout issues with OT script setup pages, that were affecting all instruments.
    • UI elements with hard-coded sizes were not getting drawn (too big to fit vertically).
    • Replaced with dynamic layouts as in rest of OT
    • Ready for team test now, svn rev 2730

KS

Last week
  • OAC HK upgrade - checked this version into SVN, got warnings coming out and setthresholds client for both alarms and warnings.
  • Meeting with Fernando and MB for LBC controller upgrade; sent Fernando the last set of config files we used while testing the BBB.
  • Installed a units update to OVMS+ during E-time.

This week
  • Add EPICS interface to OAC HK upgrade for alarming.

MB

  • management OC, WPC, ISP#2 interview
  • meeting LBC Controller upgrade, Argos, Logging, VM/SSD
  • iteration and prototype git, rpm, Jenkins CI, rpm, yum server (ctd.)
  • issuetrak automated trending
  • aggregated support calendar
Next:
  • responsibilities and assignments draft
  • techtalk#1 git//ci/rpm/kickstart
  • IT7175 Power failure follow-up meeting
  • LN / docker

PK

  • DUT1 changes (IT7084)
  • TTL (IT3949)
    • GUIs
    • TPK calculates wrong times, still chasing this

Next:
  • IT7131 (after Jose confirmation)
  • looks like again IT6746
  • TTL IT3949

SH

  • turn on VNC monitoring for obs1 -- see what John H. says.
  • changes to Zabbix archive script to ignore errors, better report errors.
  • continue work on rm507 monitor issues.
  • upgrade Louise's machine to Windows 10.

  • Work on RPMs.
  • Collect history of commands issues by observers, categorize and process.

XZ

Last Week:
  • Working on and closed few ITs (6127,7137,6396,7139,6228,5827)
  • Fastlink fiber change for LN
  • Check the current AO power failure setting
  • Finished the MG document and give feedback to MG
  • Skype call with Martin G. about the OVMS plus SW change
  • Added/changed log informtion for TT LUT
  • Reaction some operation issue (eg., the WFS high memory, Adsec flat issue, running old SW version)
This week:
  • write the low level communication code for the PowerBack Plane.

YZ

  • Try to compile the AOSupervisor code.
  • Try to understand the current design and problems in the AO logging code.
  • Understand more about the current TCS event code.
  • Understand more about the logging needs for AO code.
  • Research suitable logging implementations (e.g., spdlog) as basis for local use (or to develop our own solution).

Action items

  • xxx
  • xxx
  • xxx

-- %USERSIG{MatthieuBec}%

Comments

-- %BUBBLESIG{XianyuZhang - 2018-03-30}%
 
Topic revision: r21 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback